Are lines L1 through points (-3,2) and (0,4) and L2 through points (2,3) and (5,1) perpendicular?
Answer by checkley71(8403)

L1 HAS A SLOPE=(4-2)/(0+3)=2/3
L2 HAS A SLOPE=(1-3)/(5-2)=-2/3
THESE SLOPES ARE NOT NEGATIVE RECIPRICALS OF EACH OTHER. THEREFORE THE L1 & L2 LINES ARE NOT PERPENDICULAR.